# How Muon’s Spectral Design Benefits Generalization: A Study on Imbalanced Data

### v1

@@ B.5 Proof of Proposition˜2
Suppose Condition 1 holds. Let the weights be initialized as $\bm{W}^{0}_{0}=e^{-\delta}\bm{Q}[{\bm{I}}_{d_{1}}\,\,\mathbf{0}_{d-d_{1}}]{\bm{V}}^{\top}$ and $\bm{W}^{1}_{0}=e^{-\delta}{\bm{U}}[{\bm{I}}_{k}\,\,\mathbf{0}_{d_{1}-k}]\bm{Q}^{\top}$, where $\bm{Q}\in\mathbb{R}^{d_{1}\times d_{1}}$ is an orthonormal matrix, $k<d_{1}<d$, and $\delta>0$ is a constant. Then, for SpecGD, at each iteration, $\bm{W}^{0}_{t}:=\bm{Q}\overline{\bm{W}}^{0}_{t}{\bm{V}}^{\top}$, $\bm{W}_{t}^{1}={\bm{U}}\overline{\bm{W}}^{1}_{t}\bm{Q}^{\top}$, where $\overline{\bm{W}}^{0}_{t},\overline{\bm{W}}^{1}_{t}$ are zero except their main diagonal along which, in the limit $\delta\rightarrow\infty$, they evolve as follows for $i\in[k]$:
@@ B.6 SpecGD Dynamics for Deep Linear Model
The following result characterizes the dynamics of SpecGD for a deep $L$-layer linear model ($L>2$).
@@
Assume $k=d$, $L>2$, and suppose Ass. 1 holds. Let weights be initialized as
@@ Appendix C Additional Results in the Linear Setting
In Fig.˜10, we compare the dynamics of signGD with GD, NGD and SpecGD. We find (from the right plot) that in contrast to (N)GD and SpecGD iterates that remain in the ${\bm{U}}{\bm{V}}^{\top}$ eigenbasis (defined as in Condition 1), SignGD iterates take some time to align to that basis. We also find that SignGD learns the spectral components in a more balanced way compared to (N)GD and gets slightly better worst-class loss early on in training. However, since SpecGD learns the spectral components at an equal rate, it attains the best worst-class and class-balanced loss early on. All optimizers eventually lead to convergence to the same solution.
Figure 10: Dynamics of the iterates (left) and loss (middle) for GD, NGD, SignGD, and SpecGD on a linear model with class imbalance (priors $p_{1}\!>\!p_{2}\!>\!p_{3}$). SpecGD learns all spectral components at the same rate, whereas (N)GD prioritizes more dominant components. SignGD also seems to learn the spectral components at the same rate; however, its iterates take time to align to the ${\bm{U}}{\bm{V}}^{\top}$ eigenbasis (as seen from the residual of projected iterates (right)). Although all converge to the same solution, SpecGD significantly outperforms in worst-class and class-balanced loss early on in training.
@@
Next, we compare the effect of relaxing Condition 1 on the dynamics of the iterates. Specifically, in the relaxed version, the joint diagonalization holds with ${\bm{S}_{{\bm{x}}{\bm{x}}}}+\bm{B}$, instead of just a diagonal matrix $\bm{S}_{{\bm{x}}{\bm{x}}}$}, such that $\lVert\bm{B}\rVert\leq\epsilon$. We note that while Condition 1 is satisfied for our data model (DM) in the population limit, only the relaxed version holds in the finite sample setting. For instance, under the setting considered in Fig.˜11, $\lVert\bm{B}\rVert\approx 0.013$. In Fig.˜12, we compare the dynamics of iterates in the finite sample setting ($\lVert\bm{B}\rVert=\epsilon$) to a baseline, where we use the ${\bm{U}}{\bm{V}}^{\top}$ eigenbasis from this setting but set $\bm{B}=\mathbf{0}$. We find that the iterates in the two settings closely follow each other.
@@ Appendix D Additional Results and Details of Experimental Settings
For our experiments with Shampoo, we rely on the distributed–shampoo reference implementation of Shi et al., (2023) and use matrix inverse stabilization parameter $\varepsilon=10^{-6}$, precondition_frequency =5=5, max_preconditioner_dim =8192=8192, unless stated otherwise. To stabilize training, we also use SGDGraftingConfig or AdamGraftingConfig for the update grafting. Here, $\beta_{1}$ governs the exponential moving average of the raw gradients, while $\beta_{2}$ governs the accumulation in the matrix preconditioners (see App.˜A). Further, $\varepsilon$ is a small diagonal ‘jitter’ that stabilises the matrix inverse, and SGD/Adam grafting simply rescales the Shampoo update so its overall magnitude matches that of a plain SGD/Adam step (with default $\beta$ parameters).
For our experiments with Muon, we use the implementation from Jordan et al., (2024). Here, $\beta$ represents the momentum parameter, and we perform one Newton-Schulz iteration in each step. By default, the Adam optimizer is used to update all vector parameters. For hyperparameter tuning, we set $\beta_{1}$ for Adam to be the same as $\beta$ for Muon. We use the default value of $\beta_{2}$ for Adam with training with Muon.
For our experiments with Adam and SGD, we use the standard implementations from the PyTorch library Paszke et al., (2019). For SGD $\beta$ refers to the momentum parameter whereas for Adam $\beta_{1}$ refers to momentum and $\beta_{2}$ is the exponential moving average parameter (see App.˜A).
Following recent works (Srećković et al.,, 2025; Marek et al.,, 2025), which show that the performance gap between SGD and adaptive optimizers like Adam is larger for larger batch sizes, we use larger batch sizes in our experiments.
@@ D.1 Colored-MNIST
Dataset. In Sec.˜1, we consider a variant of the Colored-MNIST dataset, a benchmark used commonly in the literature on spurious correlations (Arjovsky et al.,, 2020; Pezeshki et al.,, 2021). The task is to classify each digit as either <5<5 or $\geq 5$. The digits in each class are injected with a background color (red or green) that is correlated with the label and acts as a spurious feature. In our setting, the spurious correlation in the train set is $99\%$.
@@
Results. Fig.˜13 shows the set of hyperparameters we tried for different optimizers and effect on majority- and minority-group test accuracy. All results are averaged over five independent runs. We select the best hyperparameter setting based on the best average group-balanced accuracy at the end of training.
Building on our initial comparison of optimizers from the Frobenius, $\ell_{\infty}$, and spectral norm families in Fig.˜1, we present a broader comparison that includes additional optimizers in Fig.˜14.
@@
Figure 14: Comparison of optimizer test accuracy on the Colored-MNIST dataset. The top and bottom panels show performance on the majority and minority groups, respectively. We compare several families of optimizers, starting with Normalized Steepest Descent methods—NGD, SignGD, and SpecGD—which normalize the gradient using the Frobenius, $\ell_{\infty}$, and spectral norms, respectively. We then show their corresponding Normalized Mometum Descent variants: NMD, Signum, and Muon. Finally, we compare other adaptive methods, contrasting Adam (element-wise preconditioning) with Shampoo (full-matrix preconditioning), along with their non-momentum analogs, RMSProp and vShampoo, to isolate the effect of preconditioning. We observe that Muon and SpecGD significantly outperform their Frobenius (NMD, NGD) and $\ell_{\infty}$ (SignGD, Signum) counterparts in terms of test accuracy on minority groups. However, we do not observe a discernible performance gap when comparing Shampoo with Adam or vShampoo with RMSProp.
@@ D.2 CIFAR-10/100 with Class Imbalance
Dataset and Preprocessing. We construct class-imbalanced versions of CIFAR-10 and CIFAR-100 (Krizhevsky et al.,, 2009) using step imbalance with ratio $R=20$. For CIFAR-10, we sample half the classes (5) to be majority classes containing $20\times$ more samples (5000) than the rest of the (minority) classes (250). The same procedure is applied to CIFAR-100 using all 100 classes.
@@
Model Architecture and Training. We train ResNet-18 on CIFAR-10 and ResNet-50 on CIFAR-100 following standard architectures (He et al.,, 2016). All models use random initialization.
@@ D.2.1 Effect of Class Imbalance Ratio
Results. Figures 19 and 20 show test accuracy dynamics across varying imbalance ratios. On CIFAR-10, Muon’s advantage grows with imbalance severity. This gap is particularly pronounced early in training: while all optimizers perform similarly on balanced data ($R=1$), at $R=20$ Muon achieves about $20\%$ minority-class accuracy compared to $10\%$ for NMD and Signum at the end of epoch $1$. On CIFAR-100, the effect of changing $R$ is less pronounced and Muon maintains the best performance across all settings.
@@ D.2.2 Fine-tuning from ImageNet Pretraining
Results. As shown in Fig. 21, when fine-tuning from ImageNet initialization, the optimizer dynamics differ from training from scratch (Fig.˜17). Adam, Muon, and Signum achieve comparable training loss and majority-class accuracy, while NMD exhibits slower convergence. For minority-class accuracy, Adam demonstrates the best performance, followed by Muon and Signum, while NMD struggles to learn minority classes quickly.
@@ D.3 MNIST-CIFAR Dataset
D.4 Subgroup-Robustness Benchmarks
@@
To ensure our comparisons are robust, we conduct extensive hyperparameter sweeps for all optimizers, as shown for the MultiNLI dataset in Fig.˜22 and CelebA in Fig.˜23. This selection protocol confirms that the consistent superior performance of spectral optimizers like Muon and Shampoo over SGD is not an artifact of a specific hyperparameter choice.
@@ D.5 TinyStories
Dataset and Preprocessing. We use a subset of the TinyStories corpus (Eldan and Li,, 2023), randomly sampling $500$ stories from the full dataset as training set. The selected stories are tokenized using SentencePiece (Kudo and Richardson,, 2018) with unigram tokenization and vocabulary size $2000$. We sampled another $50$ stories as validation set.
@@
Training is conducted with batch size $32$ for $100$ epochs using three optimizers: SGD ($\eta=0.1$, $\beta=0.9$, weight decay $10^{-4}$), Adam ($\eta=0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $10^{-4}$), and Muon ($\eta=0.1$, $\beta=0.95$, no weight decay). Learning rates and momentum values are selected via grid search based on the lowest validation loss (see Fig.˜24).
Results. As shown in Fig.˜25 (right), Muon and Adam converge faster than SGD, with Muon showing rapid loss reduction around epoch 10. Critically, SGD exhibits a large gap between majority and minority token performance, while Muon significantly reduces this gap, achieving more balanced learning across token frequencies. Also see Fig. 9 which compares the validation set accuracy, where the effect is more pronounced.
@@ D.6 Attribute-Organism Classification
Dataset and Task. The dataset contains 12 organisms across hierarchical categories: plants (apple, orange, rose, tulip), animals (dog, wolf, cat, tiger, sparrow, eagle, salmon, trout), mammals (dog, wolf, cat, tiger), and other animals. Each organism has 13 attributes with taxonomic structure: plant attributes (“uses sunlight for food,” “grows roots”), general animal attributes (“needs oxygen,” “can move around”), mammal attributes (“stays warm-blooded”), etc. The conditional probability matrix, given attribute and predict subject, is shown in Fig.˜26. This creates a hierarchical classification challenge where the model must learn both individual attribute predictions and the underlying taxonomic relationships.
@@
Results. Fig.˜26 shows confusion matrices tracking category classification for GD and Muon. The categories are organized from coarse to fine in these matrices. Each entry indicates the number of examples in the row category being classified in the column category. The results reveal a key difference in learning dynamics: GD learns the coarse categories (Animals and Plants) much faster than fine-grained categories, showing strong diagonal entries for these broad classes early in training. In contrast, Muon learns all hierarchical levels at similar rates, achieving more balanced progress across both coarse and fine categories. This differential learning behavior directly supports our theoretical finding that spectrum-aware optimizers like Muon learn principal components at equal rates, while GD prioritizes learning dominant patterns first.
@@
D.7 Effect of Loss Re-weighting

### v2

@@ B.5 Proof of Proposition˜2
Suppose Condition 1 holds. Let the weights be initialized as $\bm{W}^{0}_{0}=e^{-\delta}\bm{Q}[{\bm{I}}_{d_{1}}\,\,\mathbf{0}_{d-d_{1}}]{\bm{V}}^{\top}$ and $\bm{W}^{1}_{0}=e^{-\delta}{\bm{U}}[{\bm{I}}_{k}\,\,\mathbf{0}_{d_{1}-k}]\bm{Q}^{\top}$, where $\bm{Q}\in\mathbb{R}^{d_{1}\times d_{1}}$ is an orthonormal matrix, $k<d_{1}<d$, and $\delta>0$ is a constant. Then, for SpecGD, at each iteration, $\bm{W}^{0}_{t}:=\bm{Q}\overline{\bm{W}}^{0}_{t}{\bm{V}}^{\top}$, $\bm{W}_{t}^{1}={\bm{U}}\overline{\bm{W}}^{1}_{t}\bm{Q}^{\top}$, where $\overline{\bm{W}}^{0}_{t},\overline{\bm{W}}^{1}_{t}$ are zero except their main diagonal along which the entries evolve as follows for $i\in[k]$:
@@ B.6 SpecGD Dynamics for Deep Linear Model
The following result characterizes the dynamics of SpecGD for a deep $L$-layer linear model ($L\geq 2$).
@@
Assume $k=d$, $L\geq 2$, and suppose Ass. 1 holds. Let weights be initialized as
@@
Further, the gap between saturation of the minority vs. majority component is $\Delta T\!:=\!\tfrac{t_{\max}-t_{\min}}{t_{\min}}\!=\!\left(\tfrac{s^{\texttt{yx}}_{1}/s^{\texttt{xx}}_{1}}{s^{\texttt{yx}}_{k}/s^{\texttt{xx}}_{k}}\right)^{1/L}\!-\!1$.
@@ Appendix C Additional Results in the Linear Setting
In Fig.˜10, we compare the dynamics of SignGD, Adam (with $\beta_{1}=0.9$, $\beta_{2}=0.999$) and Muon (with $\beta=0.9$) with GD, NGD and SpecGD. We find (from the rightmost plot) that in contrast to (N)GD and SpecGD iterates that remain in the ${\bm{U}}{\bm{V}}^{\top}$ eigenbasis (defined as in Condition 1), SignGD and Adam iterates take some time to align to that basis. We also find that SignGD and Adam learn the spectral components in a more balanced way compared to (N)GD, and get slightly better worst-class loss early on in training. However, since SpecGD learns the spectral components at an equal rate, it attains the best worst-class and class-balanced loss early on. Notably, the dynamics of Muon closely follow those of SpecGD. All optimizers eventually lead to convergence to the same solution.
Figure 10: Dynamics of the iterates (left) and loss (middle) for GD, NGD, SignGD, SpecGD, Adam and Muon on a linear model with class imbalance (priors $p_{1}\!>\!p_{2}\!>\!p_{3}$). SpecGD and Muon learn all spectral components at the same rate, whereas (N)GD prioritizes more dominant components. SignGD and Adam also seem to learn the spectral components at the same rate; however, their iterates take time to align to the ${\bm{U}}{\bm{V}}^{\top}$ eigenbasis (as seen from the residual of projected iterates (right)). Although all converge to the same solution, SpecGD and Muon significantly outperform in worst-class and class-balanced loss early on in training.
Figure 12: Comparison of the dynamics of iterates (left) and loss (middle) for SpecGD (both in the population limit with zero initialization (theory) and in the finite sample setting with random initialization (expt)) and Muon (expt) on a linear model with class imbalance (priors $p_{1}\!>\!p_{2}\!>\!p_{3}$), and norm of the difference between the iterates (SpecGD-theory vs SpecGD/Muon expt) (right). We find that the dynamics characterizes by the theory closely match the experimental behaviour.
@@
In Fig.˜12, we consider the same theoretical and experimental setting as in Fig.˜11, and compare the dynamics of SpecGD in both cases to the dynamics of Muon (with $\beta=0.9$) in the experimental setting. We use a learning rate of $0.005$ in this case. We observe that the iterate and loss dynamics for Muon closely follow those predicted theoretically for SpecGD.
@@
In the next experiment, we examine the effect of relaxing Condition 1 on the dynamics of GD and SpecGD iterates. Specifically, in the relaxed version, the joint diagonalization holds with ${\bm{S}_{{\bm{x}}{\bm{x}}}}+\bm{B}$, instead of just a diagonal matrix $\bm{S}_{{\bm{x}}{\bm{x}}}$}, such that $\lVert\bm{B}\rVert\leq\epsilon$. We note that while Condition 1 is satisfied for our data model (DM) in the population limit, only the relaxed version holds in the finite sample setting. For instance, under the setting considered in Fig.˜11, $\lVert\bm{B}\rVert\approx 0.013$. In Fig.˜13, we compare the dynamics of iterates in the finite sample setting ($\lVert\bm{B}\rVert=\epsilon$) to a baseline, where we use the ${\bm{U}}{\bm{V}}^{\top}$ eigenbasis from this setting but set $\bm{B}=\mathbf{0}$. We find that the iterates in the two settings closely follow each other.
@@ Appendix D Additional Results and Details of Experimental Settings
For our experiments with Shampoo, we rely on the distributed–shampoo reference implementation of shi2023pytorchshampoo and use matrix inverse stabilization parameter $\varepsilon=10^{-6}$, precondition_frequency =5=5, max_preconditioner_dim =8192=8192, unless stated otherwise. To stabilize training, we also use SGDGraftingConfig or AdamGraftingConfig for the update grafting. Here, $\beta_{1}$ governs the exponential moving average of the raw gradients, while $\beta_{2}$ governs the accumulation in the matrix preconditioners (see App.˜A). Further, $\varepsilon$ is a small diagonal ‘jitter’ that stabilises the matrix inverse, and SGD/Adam grafting simply rescales the Shampoo update so its overall magnitude matches that of a plain SGD/Adam step (with default $\beta$ parameters).
For our experiments with Muon, we use the implementation from jordan2024muon. Here, $\beta$ represents the momentum parameter, and we perform one Newton-Schulz iteration in each step. By default, the Adam optimizer is used to update all vector parameters. For hyperparameter tuning, we set $\beta_{1}$ for Adam to be the same as $\beta$ for Muon. We use the default value of $\beta_{2}$ for Adam with training with Muon.
For our experiments with Adam and SGD, we use the standard implementations from the PyTorch library pytorch. For SGD $\beta$ refers to the momentum parameter whereas for Adam $\beta_{1}$ refers to momentum and $\beta_{2}$ is the exponential moving average parameter (see App.˜A).
Following recent works (srećković2025batchsizeproblemrevisiting; marek2025smallbatchsizetraining), which show that the performance gap between SGD and adaptive optimizers like Adam is larger for larger batch sizes, we use larger batch sizes in our experiments.
@@ D.1 Colored-MNIST
Dataset. In Sec.˜1, we consider a variant of the Colored-MNIST dataset, a benchmark used commonly in the literature on spurious correlations (irm; gs). The task is to classify each digit as either <5<5 or $\geq 5$. The digits in each class are injected with a background color (red or green) that is correlated with the label and acts as a spurious feature. In our setting, the spurious correlation in the train set is $99\%$.
@@
Results. Fig.˜14 shows the set of hyperparameters we tried for different optimizers and effect on majority- and minority-group test accuracy. All results are averaged over five independent runs. We select the best hyperparameter setting based on the best average group-balanced accuracy at the end of training.
Building on our initial comparison of optimizers from the Frobenius, $\max$, and spectral norm families in Fig.˜1, we present a broader comparison that includes additional optimizers in Fig.˜15.
@@
Figure 15: Comparison of optimizer test accuracy on the Colored-MNIST dataset. The top and bottom panels show performance on the majority and minority groups, respectively. We compare several families of optimizers, starting with Normalized Steepest Descent methods—NGD, SignGD, and SpecGD—which normalize the gradient using the Frobenius, $\max$, and spectral norms, respectively. We then show their corresponding Normalized Mometum Descent variants: NMD, Signum, and Muon. Finally, we compare other adaptive methods, contrasting Adam (element-wise preconditioning) with Shampoo (full-matrix preconditioning), along with their non-momentum analogs, RMSProp and vShampoo, to isolate the effect of preconditioning. We observe that Muon and SpecGD significantly outperform their Frobenius (NMD, NGD) and $\max$ (SignGD, Signum) counterparts in terms of test accuracy on minority groups. However, we do not observe a discernible performance gap when comparing Shampoo with Adam or vShampoo with RMSProp.
@@ D.2 CIFAR-10/100 with Class Imbalance
Dataset and Preprocessing. We construct class-imbalanced versions of CIFAR-10 and CIFAR-100 (krizhevsky2009learning) using step imbalance with ratio $R=20$. For CIFAR-10, we sample half the classes (5) to be majority classes containing $20\times$ more samples (5000) than the rest of the (minority) classes (250). The same procedure is applied to CIFAR-100 using all 100 classes.
@@
Model Architecture and Training. We train ResNet-18 on CIFAR-10 and ResNet-50 on CIFAR-100 following standard architectures (he2016deep). All models use random initialization.
@@ D.2.1 Effect of Class Imbalance Ratio
Results. Figures 20 and 21 show test accuracy dynamics across varying imbalance ratios. On CIFAR-10, Muon’s advantage grows with imbalance severity. This gap is particularly pronounced early in training: while all optimizers perform similarly on balanced data ($R=1$), at $R=20$ Muon achieves about $20\%$ minority-class accuracy compared to $10\%$ for NMD and Signum at the end of epoch $1$. On CIFAR-100, the effect of changing $R$ is less pronounced and Muon maintains the best performance across all settings.
@@ D.2.2 Fine-tuning from ImageNet Pretraining
Results. As shown in Fig. 22, when fine-tuning from ImageNet initialization, the optimizer dynamics differ from training from scratch (Fig.˜18). Adam, Muon, and Signum achieve comparable training loss and majority-class accuracy, while NMD exhibits slower convergence. For minority-class accuracy, Adam demonstrates the best performance, followed by Muon and Signum, while NMD struggles to learn minority classes quickly.
@@ D.3 MNIST-CIFAR Dataset
D.4 Subgroup Robustness Benchmarks
@@
To ensure our comparisons are robust, we conduct extensive hyperparameter sweeps for all optimizers, as shown for the MultiNLI dataset in Fig.˜23 and CelebA in Fig.˜24. This selection protocol confirms that the consistent superior performance of spectral optimizers like Muon and Shampoo over SGD is not an artifact of a specific hyperparameter choice.
@@ D.5 TinyStories
Dataset and Preprocessing. We use a subset of the TinyStories corpus (eldan-li-2023-tinystories), randomly sampling $500$ stories from the full dataset as training set. The selected stories are tokenized using SentencePiece (kudo2018sentencepiece) with unigram tokenization and vocabulary size $2000$. We sampled another $50$ stories as validation set.
@@
Training is conducted with batch size $32$ for $100$ epochs using three optimizers: SGD ($\eta=0.1$, $\beta=0.9$, weight decay $10^{-4}$), Adam ($\eta=0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $10^{-4}$), and Muon ($\eta=0.1$, $\beta=0.95$, no weight decay). Learning rates and momentum values are selected via grid search based on the lowest validation loss (see Fig.˜25).
Results. As shown in Fig.˜26 (right), Muon and Adam converge faster than SGD, with Muon showing rapid loss reduction around epoch 10. Critically, SGD exhibits a large gap between majority and minority token performance, while Muon significantly reduces this gap, achieving more balanced learning across token frequencies. Also see Fig. 9 which compares the validation set accuracy, where the effect is more pronounced.
@@ D.6 Attribute-Organism Classification
Dataset and Task. The dataset contains 12 organisms across hierarchical categories: plants (apple, orange, rose, tulip), animals (dog, wolf, cat, tiger, sparrow, eagle, salmon, trout), mammals (dog, wolf, cat, tiger), and other animals. Each organism has 13 attributes with taxonomic structure: plant attributes (“uses sunlight for food,” “grows roots”), general animal attributes (“needs oxygen,” “can move around”), mammal attributes (“stays warm-blooded”), etc. The conditional probability matrix, given attribute and predict subject, is shown in Fig.˜27. This creates a hierarchical classification challenge where the model must learn both individual attribute predictions and the underlying taxonomic relationships.
@@
Results. Fig.˜27 shows confusion matrices tracking category classification for GD and Muon. The categories are organized from coarse to fine in these matrices. Each entry indicates the number of examples in the row category being classified in the column category. The results reveal a key difference in learning dynamics: GD learns the coarse categories (Animals and Plants) much faster than fine-grained categories, showing strong diagonal entries for these broad classes early in training. In contrast, Muon learns all hierarchical levels at similar rates, achieving more balanced progress across both coarse and fine categories. This differential learning behavior directly supports our theoretical finding that spectrum-aware optimizers like Muon learn principal components at equal rates, while GD prioritizes learning dominant patterns first.
@@
D.7 Mixture of TinyStories and Add-$k$
In this section, we investigate whether our findings on imbalanced learning extend to the level of knowledge acquisition and skill learning. We introduce a domain-imbalanced setting by constructing a composite dataset where the majority component consists of natural language narratives (TinyStories), while the minority component requires learning a structured mathematical skill.
Dataset and Task. We train on a mixture of the TinyStories dataset (majority) and a mathematical task (minority), specifically add-$k$. For TinyStories, we follow the same train/test process as in Sec.˜D.5. For add-$k$, an example sequence is of the form “Q: $x_{1}$, A: $y_{1}$, Q: $x_{2}$, A: $y_{2}$, Q: $x_{3}$, A: $y_{3}$”, where $y_{j}=x_{j}+k$. We consider two-digit natural numbers for both the inputs $x_{j}$ and the offsets $k$, where $k$ is drawn from a fixed set of values during both training and evaluation. We train using the next token prediction objective and evaluate the exact match accuracy to predict the last $y_{3}$, given the rest of the sequence.
Model Architecture and Training. We use the same Transformer architecture as in our TinyStories experiments (see Sec.˜D.5 for details). We train on a composite dataset of $20,000$ samples. This includes $18,000$ TinyStories samples (majority, with $0.9$ fraction) and $2,000$ add-$k$ arithmetic sequences (minority, with $0.1$ fraction). We evaluate on $500$ held-out samples for each domain. We train for $200$ epochs with a batch size of $256$ and compare three optimizers. SGD uses $\eta=5\times 10^{-2}$, momentum $0.95$, and weight decay $10^{-4}$. For Adam, we set $\eta=10^{-3}$, $\beta_{1}=0.9,\beta_{2}=0.999$, and weight decay $10^{-4}$. For Muon, we set $\eta=10^{-2}$ and momentum $0.95$.
Results. We compare the top-1 and top-5 accuracies on TinyStories and the exact match accuracies on add-$k$ for SGD, Adam, and Muon in Fig.˜28. We observe that for Muon, the accuracies on both datasets increase at a more similar rate compared to SGD, and it generalizes better on the minority add-$k$ task. Adam, on the other hand, achieves performance comparable to Muon, similar to some of our results on subgroup robustness datasets in Sec.˜4.
Refer to caption Refer to caption Refer to caption
Figure 28: Comparison of SGD, Adam and Muon when training a transformer on a mixture of TinyStories (majority) and add-$k$ (minority) datasets: train loss (left), top-1 and top-5 test accuracies on TinyStories (middle) and exact match test accuracy on add-$k$ (right). Muon promotes more balanced learning compared to SGD while performing comparably to Adam.
Figure 30: Comparison of GD, Adam, Shampoo and Muon when training a CNN on the Barcoded MNIST dataset from kunstner2024heavy, a variant of MNIST with heavy-tailed class imbalance. GD only drives the loss on majority classes towards 0 and makes little progress on the minority classes. In contrast, Adam, Shampoo and Muon drive the loss on both majority and minority classes towards 0.
D.8 Effect of Loss Re-weighting